A father is four times as old as his daughter. In 7 years, he will be three times as old as she is. How old is the daughter now?
Answer by sachi(548)

Let the age of the daughter be x
and the age of father be 4x
After 7 yrs
the age of the daughter is x+7
the age of the father is 4x+7
ATQ 4x+7=3(x+7)
or 4x+7=3x+21
or 4x-3x=21-7
or x=14
So the age of the daughter now is 14. (Ans)